\begin{thebibliography}{HPT18}

\bibitem[Bea77]{BeauvilleThesis}
Arnaud Beauville.
\newblock Vari\'{e}t\'{e}s de {P}rym et jacobiennes interm\'{e}diaires.
\newblock {\em Ann. Sci. \'{E}cole Norm. Sup. (4)}, 10(3):309--391, 1977.

\bibitem[Bho86]{Bhosle}
Usha~N. Bhosle.
\newblock Nets of quadrics and vector bundles on a double plane.
\newblock {\em Math. Z.}, 192(1):29--43, 1986.

\bibitem[CTS87]{CTSan}
Jean-Louis Colliot-Th\'{e}l\`ene and Jean-Jacques Sansuc.
\newblock La descente sur les vari\'{e}t\'{e}s rationnelles. {II}.
\newblock {\em Duke Math. J.}, 54(2):375--492, 1987.

\bibitem[DM98]{DebMan}
Olivier Debarre and Laurent Manivel.
\newblock Sur la vari\'et\'e des espaces lin\'eaires contenus dans une
  intersection compl\`ete.
\newblock {\em Math. Ann.}, 312(3):549--574, 1998.

\bibitem[Esn03]{Esnault}
H\'{e}l\`ene Esnault.
\newblock Varieties over a finite field with trivial {C}how group of 0-cycles
  have a rational point.
\newblock {\em Invent. Math.}, 151(1):187--191, 2003.

\bibitem[FR05]{FR}
N.~Fakhruddin and C.~S. Rajan.
\newblock Congruences for rational points on varieties over finite fields.
\newblock {\em Math. Ann.}, 333(4):797--809, 2005.

\bibitem[GHS03]{GHS}
Tom Graber, Joe Harris, and Jason Starr.
\newblock Families of rationally connected varieties.
\newblock {\em J. Amer. Math. Soc.}, 16(1):57--67, 2003.

\bibitem[Hir66]{hirzebruch}
F.~Hirzebruch.
\newblock {\em Topological methods in algebraic geometry}.
\newblock Third enlarged edition. New appendix and translation from the second
  German edition by R. L. E. Schwarzenberger, with an additional section by A.
  Borel. Die Grundlehren der Mathematischen Wissenschaften, Band 131.
  Springer-Verlag New York, Inc., New York, 1966.

\bibitem[HPT18]{HPT17}
Brendan Hassett, Alena Pirutka, and Yuri Tschinkel.
\newblock Intersections of three quadrics in {$\Bbb P^7$}.
\newblock In {\em Surveys in differential geometry 2017. {C}elebrating the 50th
  anniversary of the {J}ournal of {D}ifferential {G}eometry}, volume~22 of {\em
  Surv. Differ. Geom.}, pages 259--274. Int. Press, Somerville, MA, 2018.

\bibitem[HRS18]{HRS}
Jonathan~D. Hauenstein, Jose~Israel Rodriguez, and Frank Sottile.
\newblock Numerical computation of {G}alois groups.
\newblock {\em Found. Comput. Math.}, 18(4):867--890, 2018.

\bibitem[HT10]{HT10}
Brendan Hassett and Yuri Tschinkel.
\newblock Flops on holomorphic symplectic fourfolds and determinantal cubic
  hypersurfaces.
\newblock {\em J. Inst. Math. Jussieu}, 9(1):125--153, 2010.

\bibitem[Jia12]{Jiang}
Zhi Jiang.
\newblock A {N}oether-{L}efschetz theorem for varieties of {$r$}-planes in
  complete intersections.
\newblock {\em Nagoya Math. J.}, 206:39--66, 2012.

\bibitem[Las89]{Laszlo}
Yves Laszlo.
\newblock Th\'{e}or\`eme de {T}orelli g\'{e}n\'{e}rique pour les intersections
  compl\`etes de trois quadriques de dimension paire.
\newblock {\em Invent. Math.}, 98(2):247--264, 1989.

\bibitem[Man91]{Manivel}
Laurent Manivel.
\newblock Un th\'eor\`eme d'annulation pour les puissances ext\'erieures d'un
  fibr\'e ample.
\newblock {\em J. Reine Angew. Math.}, 422:91--116, 1991.

\bibitem[O'G86]{ogrady86}
Kieran~G. O'Grady.
\newblock The {H}odge structure of the intersection of three quadrics in an
  odd-dimensional projective space.
\newblock {\em Math. Ann.}, 273(2):277--285, 1986.

\bibitem[Shi04]{Shimada}
Ichiro Shimada.
\newblock Vanishing cycles, the generalized {H}odge conjecture and
  {G}r\"{o}bner bases.
\newblock In {\em Geometric singularity theory}, volume~65 of {\em Banach
  Center Publ.}, pages 227--259. Polish Acad. Sci. Inst. Math., Warsaw, 2004.

\bibitem[Tju75]{Tjurin}
A.~N. Tjurin.
\newblock The intersection of quadrics.
\newblock {\em Uspehi Mat. Nauk}, 30(6(186)):51--99, 1975.

\bibitem[Voi10]{VoisinConiveau}
Claire Voisin.
\newblock Coniveau 2 complete intersections and effective cones.
\newblock {\em Geom. Funct. Anal.}, 19(5):1494--1513, 2010.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath,amssymb,xcolor}

\title{Varieties of planes on intersections of three quadrics}
\author{Brendan Hassett and Yuri Tschinkel}
\date{April 2019}

\theoremstyle{plain}

\theoremstyle{plain}
\newtheorem{prop}{Proposition}

\newtheorem{theo}[prop]{Theorem}

\theoremstyle{definition}

\newtheorem{ques}[prop]{Question}

\newtheorem{rema}[prop]{Remark}

\newcommand{\bC}{\mathbb C}

\newcommand{\bG}{\mathbb G}
\newcommand{\bP}{\mathbb P}
\newcommand{\bQ}{\mathbb Q}
\newcommand{\bZ}{\mathbb Z}

\newcommand{\cC}{\mathcal C}
\newcommand{\cE}{\mathcal E}
\newcommand{\cI}{\mathcal I}
\newcommand{\cO}{\mathcal O}

\newcommand{\cX}{\mathcal X}

\newcommand{\ra}{\rightarrow}

\newcommand{\Bl}{\operatorname{Bl}}
\newcommand{\Br}{\operatorname{Br}}
\newcommand{\Gr}{\operatorname{Gr}}

\newcommand{\Sym}{\operatorname{Sym}}
\newcommand{\Pic}{\operatorname{Pic}}
\newcommand{\prim}{\operatorname{prim}}

\begin{document}
\maketitle

\begin{abstract}
We study the geometry of spaces of planes on smooth complete intersections of three quadrics, with a view toward rationality questions.
\end{abstract}

\section{Introduction}

This note studies the geometry of smooth complete intersections of three quadrics $X\subset \bP^n$, with a view toward rationality questions over nonclosed fields $k$.  We review what is known over $\bC$:
\begin{itemize}
\item{$X$ is irrational for $n\le 6$ \cite{BeauvilleThesis};}
\item{$X$ may be either rational or irrational for $n=7$, and the rational ones are dense in moduli \cite{HPT17};}
\item{$X$ is always rational for $n\ge 8$ \cite[Cor.~5.1]{Tjurin}.}
\end{itemize}
The analysis in higher dimensions relies on the geometry of planes in $X$. 
Indeed, when $X$ contains a plane $P$ defined over $k$ then projection from $P$ gives a birational map
$$\pi_{P}: X \stackrel{\sim}{\dashrightarrow} \bP^{n-3}.$$
This leads us to study the variety of planes $F_2(X) \subset \Gr(3,n+1)$. When $n\ge 12$, the geometry
of these varieties gives a quick and uniform proof of rationality over finite fields and function fields 
of complex curves (see Theorem~\ref{theo:high}).  

We are therefore interested in the intermediate cases $n=8,9,10,11$, and especially in $n=8$ and $9$. 
For generic $X\subset \bP^8$, the variety $F_2(X)$ is finite of degree $1024$ (see Proposition~\ref{prop:finite});
we explore the geometry of the associated configurations of planes in $\bP^8$. We then focus most on the
case $n=9$. Here the variety $F_2(X)$ is a threefold of general type with complicated geometry -- we analyze its numerical invariants.

Our original motivation was to understand certain {\em singular} complete intersections of three quadrics in 
$\bP^9$ associated with universal torsors over degree 4 del Pezzo surfaces fibered in conics over $\bP^1$ \cite{CTSan}.
Conjectures of Colliot-Th\'el\`ene and Sansuc predict that such torsors are {\em rational} when they admit a point, over number fields.
Rationality of torsors has significant arithmetic applications, e.g., to proving the uniqueness of the Brauer--Manin obstruction
to the Hasse principle and weak approximation. It has geometric consequences as well, e.g., 
the construction of new examples of nonrational but stably rational threefolds over $\bC$. 
The geometry of {\em smooth} intersections, presented here, turned out to be quite rich and interesting on its own.    

Here is a road map of the paper: 
Section~\ref{sect:high} presents uniform proofs of rationality for high-dimensional cases. Section~\ref{sect:determinant} is
devoted to determinantal presentations of plane curves arising as degeneracy loci of the net of quadrics. 
In Sections~\ref{sect:schubert} and \ref{sect:koszul}, we turn to numerical invariants of the variety of planes. 
Our main results concern the computation of 
degrees and 
cohomology 
of $F_2(X)$, for $X$ a smooth intersection of three quadrics, in $\bP^8$ and $\bP^9$. 

\

\noindent
{\bf Acknowledgments:} 
The first author was partially supported by NSF grants 1551514 and 1701659, and the Simons Foundation; the second author was partially supported by NSF grant 1601912.

\section{Uniform rationality in high dimensions}
\label{sect:high}
\begin{theo}  \label{theo:high}
Let $k$ be a finite field or the function field of a complex curve. 
Suppose that $X\subset \bP^n, n\ge 12,$ is a smooth complete
intersection of three quadrics. Then $X$ is rational.
\end{theo}
\begin{proof}
The results of \cite{DebMan} show that for $X$ generic over $\bar{k}$,
the variety $F_2(X)\subset \Gr(3,n+1)$ is smooth and connected of the expected dimension $3n-24$. 
The adjunction formula implies that $F_2(X)$ has canonical class
$$K_{F_2(X)} = (12 - n-1) \sigma_1;$$
here $\sigma_1$ is the hyperplane class from the Pl\"ucker embedding -- see Section~\ref{sect:schubert}
for details. In particular, $F_2(X)$ is Fano for $n\ge 12$. 

Suppose $k$ is a function field. When $F_2(X)$ is smooth of the expected dimension, we have
$F_2(X)(k)\neq \emptyset$, by the Graber-Harris-Starr theorem \cite{GHS}. 
What if $F_2(X)$ is singular or fails to have the expected dimension?
However, we still have rational points, by an argument of Starr: Suppose that $\cX_{\circ} \ra B$ is a family 
of varieties over a complex projective curve $B$ corresponding to a morphism $B \ra \operatorname{Hilb}$
to the appropriate Hilbert scheme. There exists a one-parameter family of curves $\cC_t$, meeting
the locus over which $F_2$ is smooth of the expected dimension, such that $\cC_0$ contains $B$
as an irreducible component. (Take the $\cC_t$ to be complete intersection curves in $\operatorname{Hilb}$.)
For the induced families $\cX_t \ra \cC_t$, the fibrations $F_2(\cX_t)\ra \cC_t$ admit sections
for most $t$. This remains true as $t\ra 0$; restricting to $B$ gives a section of $F_2(\cX_{\circ})\ra B$.

A similar argument holds over finite fields, using Esnault's Theorem \cite{Esnault} and the specialization
version due to Fakhruddin and Rajan \cite{FR}.
\end{proof}

\section{Determinantal representations}
\label{sect:determinant}

\subsection{Recollection of invariants} \label{subsect:recall}
We recall formulas that may be obtained from Appendix I of
Hirzebruch's {\em Topological Methods}, specifically,
\cite[Th.~22.1.1]{hirzebruch}.

\begin{prop}
Suppose that $n=2m$. Then we have
$$\chi(X)=-4m(m-1)$$
and
$$h^{m-1,m-2}(X)=h^{m-2,m-1}(X)=\binom{2m}{2}-1=2m^2-m-1.$$
The other Hodge numbers of weight $2m-3$ vanish.
\end{prop}

\begin{prop}
Suppose that $n=2m-1$. Then we have
$$\chi(X)=4m(m-1)$$
and
$$h^{m-1,m-3}(X)=h^{m-3,m-1}(X)=\binom{m-1}{2}, 
h^{m-2,m-2}(X)=3m^2-3m+2.$$
\end{prop}
The odd case follows from \cite{ogrady86} and \cite{Laszlo}; the even case
from \cite{BeauvilleThesis}. For reference, we give a tabulation of nontrivial middle Hodge numbers
$$
\begin{array}{r|c}
n & \text{Hodge numbers}\\
\hline
4 & 5 \quad 5 \\
5 & 1 \quad 20 \quad 1 \\
6 & 14 \quad 14 \\
7 & 3 \quad 38 \quad 3 \\
8 & 27 \quad 27 \\
9 & 6 \quad 62 \quad 6 \\
10 & 44 \quad 44 
\end{array}
$$

\subsection{Interpretation of cohomology}
Let $X\subset \bP^n$ be a smooth complete intersection of three
quadrics containing a plane $P$. It follows that $n\ge 7$ and $X$ is of
special moduli when $n=7$, as the expected dimension of the 
Fano variety of planes is
$$\operatorname{expdim} F_2(X) = 3(n-2) - 3\cdot 6 = 3n-24.$$
This case is studied in depth in \cite{HPT17}.

Projection from $P$ induces a birational map 
$X \stackrel{\sim}{\dashrightarrow}\bP^{n-3}$. The center $Y$ of the inverse has 
dimension $n-5$ and admits a fibration in quadric hypersurfaces of relative
dimension $n-7$ over $\bP^2$.
This is instrumental in computing and interpreting the cohomology of $X$.

Suppose $n$ is odd. Then $Y$ is fibered in even-dimensional quadrics and its
cohomology is governed by the associated double cover 
$$
S \ra \bP^2,
$$
branched over the degeneracy locus $D$, and the associated Brauer class $\eta \in \Br(S)[2]$,
when $n\ge 9$. It is possible to interpret $F_2(X)$ via moduli spaces of vector bundles over
$S$  \cite{Bhosle}.

When $n$ is even, $Y$ is fibered in odd-dimensional quadrics, degenerate over a plane
curve $D$ of degree $n+1$. The cohomology is obtained via a Prym construction arising from 
the associated double cover of $D$. 

\subsection{Equations of the center}

In general, write
$$\cE:= \cO^3_{\bP^{n-3}} \oplus \cO_{\bP^{n-3}}(-1) \hookrightarrow 
\cO_{\bP^{n-3}}^{n+1}$$
so that $\bP(\cE)=\Bl_P(\bP^n)$ with bundle morphism 
$\varpi:\bP(\cE) \ra \bP^{n-3}$. 
Quadrics containing $P$ correspond to elements
of 
$$\Gamma(\cO_{\bP(\cE)}(1) \otimes \varpi^*\cO_{\bP^{n-3}}(1)) = 
\Gamma(\cE^*(1)),
$$
and complete intersections of three such forms correspond to linear
transformations
$$\cO_{\bP^{n-3}}^3 \ra \cE^*(1)\simeq \cO_{\bP^{n-3}}(1)^3 \oplus \cO_{\bP^{n-3}}(2).$$
We are interested in their degeneracy loci, whose equations are given 
by minors of            
\begin{equation} \label{eqn:4by3}
\left( \begin{matrix} L_{11} & L_{12} & L_{13} \\
                        L_{21} & L_{22} & L_{23} \\
                        L_{31} & L_{32} & L_{33} \\
                        Q_1 & Q_2 & Q_3 \end{matrix} \right) 
\end{equation}
where the $L_{ij}$ are linear and the $Q_i$ are quadratic; maximal
minors yield one cubic equation and three quartic equations.

Let $Y\subset \bP^{n-3}$ denote the subscheme given by the maximal minors,
which generically has codimension two. The $(n+1)$-dimensional
linear series
$\Gamma(\cI_Y(4))$
induces a birational map
$$\bP^{n-3} \dashrightarrow X \subset \bP^n$$
inverse to $\pi_P$.  

The determinantal equations (\ref{eqn:4by3}) for $Y$ yield a rational map
$$\phi: Y \dashrightarrow \bP^2,$$
assigning to each rank-two matrix its one-dimensional kernel.
This fails to be defined along the locus $Z \subset Y$ where the
matrix (\ref{eqn:4by3}) has rank one. See Section~\ref{subsect:generic}
for more information about its geometry.

We shall analyze the cohomology of $X$ using the
cohomology of natural resolutions of $Y$, expressed via $\phi$ as quadric
bundles over $\bP^2$.

\subsection{Linear algebra}
We start with some linear algebra on generic $3\times 3$
matrices
$$R_1:=\bP^2 \times \bP^2  \subset R_2 \subset \bP^8,$$
defined by the various minors of $L=(L_{ij})$. 
Over $R_1$, we may write 
$$L_{ij}=u_iv_j, \quad i,j=1,2,3.$$
The $2\times 2$ minors induce the `Cramer map'
$$\iota: \Bl_{R_1}(\bP^8) \ra \bP^8.$$
Let $E$ denote the exceptional divisor and $\widetilde{R_2}$
the proper transform of $R_2$.
The target space is also stratified by rank
$$R'_1 \subset R'_2 \subset \bP^8$$
with $\iota(\widetilde{R_2}) \subset R'_1$ and $\iota(E)=R'_2.$

Let $M=(M_{ij})$ denote the natural coordinates on the target $\bP^8$,
so the graph of $\iota$ is given by
$$L\cdot M = M\cdot L = \text{diagonal matrix}.$$ 
In other words, we have the equations
$$ \sum_j L_{ij}M_{jk} = \sum_{\ell} M_{i\ell} L_{\ell k}=0 \quad i\neq k$$
and 
$$ \sum_j L_{ij}M_{ji} = \sum_{\ell} M_{k\ell} L_{\ell k} \quad i=k.$$
Thus $\widetilde{R_2}$ is given by
$$\det(L_{ij})=0, \quad M_{ij}M_{k\ell}-M_{i\ell}M_{kj}=0, \quad
L\cdot M = M\cdot L =0.$$
Writing $M_{ij}=x_iy_j$ yields equations
$$
\sum_j L_{ij}x_j = \sum_{\ell} y_{\ell} L_{\ell k}=0,
$$
for all indices $i,k$.

We obtain two small resolutions of $R_2$ by keeping track of 
just the kernel or the image of $L$, respectively. For example, we
may consider
$$\widehat{R_2} \subset \bP^8_{L_{ij}} \times \bP^2_{x_j}$$
given by the equations
$$L \left( \begin{matrix} x_1 \\ x_2 \\ x_3 \end{matrix} \right)=0.$$
The exceptional locus $\widehat{E} \subset \widehat{R_2}$ is a
$\bP^1$-bundle over $R_1$ -- for each rank one matrix we extract the
one-dimensional subspaces of its kernel. 
It has codimension two in $\widehat{R_2}$.

\subsection{Generic behavior} \label{subsect:generic}
Assume that $n=11$ and the $L_{ij}$ are
linearly independent.

On the locus $R_1$ where the upper $3\times 3$ matrix
has rank at most one, we 
may write $L_{ij}=u_iv_j$.
The remaining equations involving the $Q_i$ take the form
$$u_1v_2Q_1-u_1v_1Q_2 = u_2v_1Q_1-u_2v_2Q_2 = \cdots = 0.$$
Writing $q_i=Q_i(u_1v_1,\ldots,u_3v_3)$, we obtain
$$v_2q_1 - v_1q_2 = v_3q_1 - v_1q_3 = v_2 q_3 - v_3 q_2 =0.$$
Thus there exists an
$$F \in k[u_1,u_2,u_3;v_1,v_2,v_3]_{(2,1)}$$
such that 
$$
q_1=v_1F, \quad q_2=v_2F, \quad q_3=v_3F.
$$
The indeterminacy locus for the kernel map
$\phi$ is given by a bidegree $(2,1)$ hypersurface
$$
Z \subset \bP^2 \times \bP^2.
$$  

Resolving the kernel map
$\phi: Y \dashrightarrow \bP^2$ yields a 
small resolution $\widehat{Y} \ra Y$ with center $Z$.
Write the kernel of (\ref{eqn:4by3}) as 
$[x_1,x_2,x_3]$
so that
$$[L_{12}L_{23}-L_{13}L_{23},-L_{11}L_{23}+L_{13}L_{23},L_{11}L_{22}-L_{12}L_{21}]\sim[x_1,x_2,x_3]$$
etc.
The rank two matrices with this kernel correspond to a codimension-three
linear subspace 
$$\Lambda_{[x_1,x_2,x_3]}\simeq \bP^5 \subset \bP^8.$$
The $3\times 3$ minors involving the $Q_i$ are all proportional to
$$Q_1x_1+Q_2x_2+Q_3x_3.$$
Hence 
$$\phi:\widehat{Y} \ra \bP^2$$
is a quadric bundle of relative dimension four.

\subsection{Degenerate cases in small dimension}

\

\noindent
{$\mathbf{n=5}$}: 
Suppose we have a complete intersection of three quadrics 
in $\bP^5$ containing a plane
$$X=P \cup_B U.$$
Here $\pi_P:U \ra \bP^2$ is birational, realizing
$U=\Bl_Y(\bP^2)$, where 
$$\phi:Y=\{y_1,\ldots,y_9\} \hookrightarrow \bP^2$$
is a generic collection of nine points, and the imbedding
$$
Y \hookrightarrow \bP^5
$$ 
is via quartics vanishing at those nine points. 

Note there is a canonical cubic curve
$W \supset Y$ where the determinant of linear forms
vanishes. For each divisor
$$\Sigma \equiv_W 4g-Y$$
there is a unique quartic form -- modulo the defining equation of $W$ --
cutting out $\Sigma \cup Y$. 

\

\noindent
{$\mathbf{n=6}$}: 
Now consider a complete intersection of three quadrics in $\bP^6$ containing
a plane
$$
P \subset X \subset \bP^6.
$$
The threefold $X$ is a nodal Fano threefold, with six singularities along 
$P$. Note that $X$ depends on $21$ parameters, codimension six in
the parameter space of all complete intersections.

In this case, we have
$$
Y \subset W \subset \bP^3,
$$
where $W$ is a smooth cubic surface, $Y$ is a smooth curve with
$$
\deg(Y)=9, \quad \operatorname{genus}(Y)=9,
$$
residual to a twisted cubic $\Sigma \subset W$ in the complete intersection
of $W$ and a quartic.  
Write
$$\Pic(W)=\left<L,E_1,E_2,E_3,E_4,E_5,E_6\right>, \quad L=[\Sigma], $$
where the $E_i$ are pairwise disjoint $(-1)$-classes with $L\cdot E_i=0$.
Then we have
$$[Y]=11L-4E_1-\ldots-4E_6$$
and the residual twisted cubic to $\Sigma$, 
with divisor class
$$
5L-2E_1-\ldots-2E_6,
$$ 
realizes $Y$ is a septic plane curve with
six nodes. The corresponding linear series induces $\phi:Y \ra \bP^2$. 

The intermediate Jacobian of a minimal resolution
$\widetilde{X} \ra X$ is isomorphic to the Jacobian of $Y$.

\

\noindent
{$\mathbf{n=7}$}: 
The simplest smooth case, obtained by
taking a codimension-four linear section of the generic case.   
For generic choices of the linear section, $Z=\emptyset$ and
$\phi:Y \ra \bP^2$ is a double cover branched over a special octic
plane curve.
Thus we have 
$$
Y \subset W:=\{\det(L_{ij})=0 \} \subset \bP^4,
$$
where $W$ is a cubic threefold with six ordinary double points \cite{HT10}.

The surface $Y$ has Picard group
$$\begin{array}{c|cc}
    & K & h \\
\hline
K  &  2 & 7 \\
h  &  7 & 9
\end{array}
$$
where $K$ is the canonical class -- pulled back from $\bP^2$ -- and
$h$ is associated with the embedding 
$$Y\hookrightarrow W \hookrightarrow \bP^4.$$
There is a second morphism $Y \rightarrow \bP^2$ associated with the
divisor $2h-K$ because
\begin{align*}
\chi(2h-K)& =\frac{(2h-K)^2-(2h-K)K}{2}+4 \\
                &= \frac{36-28+2-14+2}{2}+4=3.
\end{align*}
The residual intersection to $Y$ in the complete intersection of
$W$ and a quartic hypersurface is a cubic scroll $\Sigma$.
By \cite{HT10}, $Y$ admits two families of such scrolls,
each parametrized by $\bP^2$. An adjunction computation shows that
$$\Sigma \cap W \equiv_Y 2h-K.$$  

\begin{rema}
The lattice $\left<K,h\right>$ has discriminant $-31$. 
Consider the lattice
$$\left<g^2,P\right> \subset H^4(X,\bZ)$$
under the intersection pairing, where $g$ is the hyperplane class.
A Chern-class computation gives 
$P^2=4$ whence
$$\begin{array}{c|cc}
    & g^2 & P \\
\hline
g^2  &  8 & 1 \\
P    &  1 & 4
\end{array},
$$
which has discriminant $31$. 
The birational parametrization of $X$ induces an isomorphism of
Hodge structures
$$H^2(Y,\bZ)(-1) \supset \left< K,h\right>^{\perp} \simeq
\left< g^2,P\right>^{\perp} \subset H^4(X,\bZ).$$
\end{rema}

\

\noindent
{$\mathbf{n=8}$}: 
Here the $L_{ij}$ are linear forms on $\bP^5$ so the determinant cubic
is singular along an elliptic normal curve.  The center $Y\subset \bP^5$ of the mapping to $\bP^8$
is birational to a conic bundle over $\bP^2$. 
The locus $Z\subset Y$, where $Y$ meets the elliptic normal curve, 
consists of nine points.
The equations of $\widehat{Y} \subset \bP^5\times \bP^2$ may be written:
$$L_{i1}x_1+L_{i2}x_2+L_{i3}x_3=0, \quad
Q_1x_1 + Q_2x_2 + Q_3x_3=0.$$
The linear forms induce
$$0 \ra K \ra \cO_{\bP^2}^{\oplus 6} \ra \cO_{\bP^2}(1)^{\oplus 3} \ra 0 $$
and the quadratic form induces a symmetric map
$$K \ra K^*(1)$$
with degeneracy $C\subset \bP^2$ of degree $9$.  

We count parameters: the linear terms depend on $45$ parameters, the quadratic
on an additional $44$ parameters, and coordinate changes account for $43$ parameters,
leaving a total of $46$ parameters. We know that $C$ comes with a distinguished
double cover $\widetilde{C}\ra C$ but even after fixing this we obtain a total of $1024$ determinantal representations (see Proposition~\ref{prop:finite}).

\

\noindent
{$\mathbf{n=9}$}: 
Here the $L_{ij}$ are linear forms on 
$\bP^6$. This allows us to use the normal form
$$\left( \begin{matrix} A_1 & A_2 & A_3 \\
			A_5 & A_1 & A_4 \\
			A_6 & A_7 & A_1  \end{matrix} \right)
\left( \begin{matrix} x_1 \\ x_2 \\ x_3 \end{matrix} \right)=0.$$
The rank-one locus is a degree-six del Pezzo surface.
The locus $Z\subset Y$ is a curve of genus three and degree nine.
The equations of $\widehat{Y} \subset \bP^6\times \bP^2$ may be written:
\begin{align*}
A_1x_1+A_2x_2+A_3x_3=A_5x_1+A_1x_2+A_4x_3=A_6x_1+A_7x_2+A_1x_3 =&0 \\
Q_1x_1 + Q_2x_2 + Q_3x_3=&0.
\end{align*}
The linear forms induce
$$0 \ra K \ra \cO_{\bP^2}^{\oplus 7} \ra \cO_{\bP^2}(1)^{\oplus 3} \ra 0 $$
and the quadratic form induces a symmetric map
$$K \ra K^*(1)$$
with degeneracy $C\subset \bP^2$ of degree $10$.  

Fixing $C$ and the double cover $\widetilde{C}\ra C$, the various
determinant representations as above are parametrized by the
planes in $X$, i.e., by the threefold $F_2(X)$.

\section{Schubert calculus of the variety of planes}
\label{sect:schubert}

Now assume that the variety of planes
$$F_2(X) \subset \Gr(3,n+1)$$
is smooth of the expected dimension $3n-24$. 

Consider the canonical exact sequence over $\Gr(3,n+1)$
$$0 \ra S \ra V\otimes \cO \ra Q \ra 0.$$
The defining equations
$$
X = \{F_1=F_2=F_3=0 \}, \quad F_1,F_2,F_3 \in \Sym^2(V^*),
$$
induce sections $f_1,f_2,f_3 \in \Sym^2(S^*)$ so that
$$F_2(X) = \{ f_1=f_2=f_3=0\}.$$

We compute Chern classes. The Chern classes of $S$ are the Schubert
cycles
$$c(S) = 1 - \sigma_{1} + \sigma_{1,1} - \sigma_{1,1,1},
$$
yielding
$$c(S^*) = 1 + \sigma_{1} + \sigma_{1,1} + \sigma_{1,1,1}.$$
A computation in symmetric functions yields
\begin{eqnarray*}
c_1(\Sym^2(S^*)) &=& 4\sigma_{1} \\
c_2(\Sym^2(S^*)) &=& 5\sigma^2_{1}+5\sigma_{1,1}=5\sigma_{2,0}+10\sigma_{1,1} \\
c_3(\Sym^2(S^*)) &=& 2\sigma_{1}^3 + 11 \sigma_{1}\sigma_{1,1}+7\sigma_{1,1,1} \\
		 &=& 2\sigma_3 + 15\sigma_{2,1} + 20\sigma_{1,1,1} \\
c_4(\Sym^2(S^*)) &=& 6\sigma^2_{1}\sigma_{1,1} + 14 \sigma_{1}\sigma_{1,1,1}
+ 4 \sigma_{11}^2 \\
		 &=& 6\sigma_{3,1} + 10\sigma_{2,2} + 30\sigma_{2,1,1} \\
c_5(\Sym^2(S^*)) &=& 8\sigma^2_{1}\sigma_{1,1,1}+ 4\sigma_{1}\sigma_{1,1}^2
+4\sigma_{1,1}\sigma_{1,1,1} \\
		 &=& 4\sigma_{3,2} + 12 \sigma_{3,1,1} + 20\sigma_{2,2,1} \\
c_6(\Sym^2(S^*)) &=& 8 \sigma_{1}\sigma_{1,1}\sigma_{1,1,1}-8\sigma_{1,1,1}^2
			= 8 \sigma_{3,2,1}.
\end{eqnarray*}
Computing via the `SchurRings' package of \texttt{Macaulay2}, we obtain
{\small
$$[F_2(X)]=512(\sigma_{9,6,3}+2 \sigma_{9,5,4} + 2\sigma_{8,7,3}+6\sigma_{8,6,4}
+4\sigma_{8,5,5} + 4\sigma_{7,7,4}+8\sigma_{7,6,5}+2\sigma_{6,6,6}).$$
}
The paper \cite{DebMan} explores these formulas more systematically;
our formula extends the tabulation on \cite[p.~563]{DebMan}.
See also \cite{Jiang} for more information on Noether-Lefschetz
questions for Fano schemes.

\begin{prop}
\label{prop:finite}
When $n=8$ and $X$ generic, the variety  $F_2(X) \subset \Gr(3,9)$ has
dimension zero and 
$$[F_2(X)]=1024[\operatorname{point}].$$
\end{prop}

\begin{ques}
Describe the Galois action arising in this case.
To what extent is it governed by the Galois representation
on the intermediate Jacobian?
\end{ques}
Frank Sottile and his collaborators (see, e.g., \cite{HRS}) are developing computational 
approaches to Galois groups of enumerative questions. Computations
with Taylor Brysiewicz indicate the group in this case might be smaller
than the symmetric group $\mathfrak{S}_{1024}$ but a full computation
appears difficult with existing techniques and computational resources.
\begin{rema}
We propose to construct these examples synthetically: Fix seven generic planes
$$P_1,\ldots,P_7 \subset \bP^8 $$
which depend up to projectivity on $18\cdot 7 - 80 = 46$ parameters. 
Write 
$$\Pi=P_1 \sqcup P_2 \sqcup \cdots \sqcup P_7$$
and note that
$$h^0(\cI_{\Pi}(2))=h^0(\cO_{\bP^8}(2)) - 7\cdot 6 = 3.$$
Thus there is a {\em unique} complete intersection of three quadrics
$X\supset \Pi$ which we expect is smooth. How do we construct the 
other $1017$ planes of $F_2(X)$?
\end{rema}

We focus next on the case $n=9$ where $F_2(X) \subset \Gr(3,10)$
is a threefold with class
$$[F_2(X)]=512 (4\sigma_{7,7,4}+8\sigma_{7,6,5}+2\sigma_{6,6,6}).$$
We have
$$\sigma_1^3 = \sigma_3 + 2\sigma_{2,1}+\sigma_{1,1,1}$$
hence we have
$$\deg(F_2(X))=11264 = 11\cdot 2^{10}.$$
The adjunction formula implies
$$K_{F_2(X)}=2\sigma_1$$
whence
$$K_{F_2(X)}^3=11 \cdot 2^{13}.$$

\begin{prop} \cite[Th.~3.4]{DebMan} \label{prop:DebMan}
When $F_2(X)$ is smooth of the expected dimension we have
$$h^{0,1}(F_2(X))=h^{1,0}(F_2(X))=0.$$
\end{prop}

Consider the rank $18$ vector bundle
$$\cE = \Sym^2(S^*)^{\oplus 3}$$
so there exists a section $s\in \Gamma(\Gr(3,10),\cE)$
with $F_2(X)=\{s=0\}$. 
We have
$$\begin{array}{rcl}
c_1(\cE) &=& 3c_1(\Sym^2(S^*))=12\sigma_1 \\
c_2(\cE) &=& 3c_1(\Sym^2(S^*))^2+ 3c_2(\Sym^2(S^*))= 63 \sigma_2+78\sigma_{1,1}\\
c_3(\cE) &=& 190\sigma_3 + 533 \sigma_{2,1} + 364 \sigma_{1,1,1}. 
\end{array}
$$

The exact sequence
$$ 0 \ra T_{F_2(X)} \ra T_{\Gr(3,10)}|F_2(X) \ra N_{F_2(X)/\Gr(3,10)} \ra 0$$
and the interpretation of the normal bundle as $\cE|F_2(X)$
allows us to compute all the Chern classes of $T_{F_2(X)}$.

Indeed, $T_{\Gr(3,10)}=\operatorname{Hom}(S,Q)$ where 
$$c_1(Q)=\sigma_1, \quad c_2(Q)=\sigma_2, \quad c_3(Q)=\sigma_3.$$
$$
\begin{array}{rcl}
c_1(T_{\Gr(3,10)}) &=& r(S) c_1(Q) - r(Q) c_1(S) = 10\sigma_1 \\
c_2(T_{\Gr(3,10)}) &=& \binom{r(S)}{2}c_1(Q)^2 + r(S) c_2(Q) +  \binom{r(Q)}{2}c_1(S)^2 \\
& & 
		      \quad\quad\,   + r(Q) c_2(S) 
			- (r(S)r(Q)-1)c_1(Q)c_1(S) \\
		  &=& 3 \sigma_1^2 + 3 \sigma_2 + 21 \sigma_1^2 +
			7 \sigma_{1,1} -20 \sigma_1^2 \\
		  &=& 47 \sigma_2 + 51 \sigma_{1,1} \\
c_3(T_{\Gr(3,10)}) &=& 140 \sigma_3 + 310\sigma_{2,1} + 180\sigma_{1,1,1}.
\end{array}
$$
The last is computed using the identity
$$
\operatorname{ch}(T_{\Gr(3,10)})=\operatorname{ch}(S^*)\operatorname{ch}(Q),
$$
which yields
$$\operatorname{ch}_2(T_{\Gr(3,10)}) = 3\sigma_{2}-\sigma_{11}$$
and
$$\operatorname{ch}_3(T_{\Gr(3,10)}) = \frac{5}{3}(\sigma_{3}-\sigma_{2,1}+
					\sigma_{1,1,1}).$$
Note that
$$\operatorname{ch}_3(V)= \frac{1}{6}(c_1(V)^3-3c_1(V)c_2(V)+3c_3(V)).$$

We compute the Chern classes, identifying bundles with their
restrictions to $F_2(X)$:
As we have seen
$$c_1(T_{F_2(X)})= c_1(T_{\Gr(3,10)}) - c_1(\cE) = -2\sigma_1.$$
The second Chern class is:
$$
\begin{array}{rcl}
c_2(T_{F_2(X)}) &=& c_2(T_{\Gr(3,10)})-c_2(\cE)
			-c_1(T_{F_2(X)})c_1(\cE) \\
		&=& 8\sigma_2 - 3\sigma_{1,1}
\end{array}
$$
And finally
$$
\begin{array}{rcl}
c_3(T_{F_2(X)}) &=& c_3(T_{\Gr(3,10)})-c_2(T_{F_2(X)})c_1(\cE)
			-c_1(T_{F_2(X)})c_2(\cE) - c_3(\cE) \\
		&=& -20\sigma_3  - \sigma_{2,1}  + 8\sigma_{1,1,1}
\end{array}
$$

A computation with
the `SchurRings' package of \texttt{Macaulay2} gives
$$\chi(\cO_{F_2(X)})=c_1(T_{F_2(X)})c_2(T_{F_2(X)})/24 = -2816=-2^7\cdot 11$$
and
$$\chi(F_2(X))=c_3(T_{F_2(X)}) = -36,864 = -2^{12}\cdot 3^2 .$$

The Hirzebruch-Riemann-Roch formula gives
$$\chi(\cO_{F_2(X)}(1)) = 0,$$
which is to be expected, as Serre duality gives
$$h^i(\cO_{F_2(X)}(1))=h^{3-i}(\omega_{F_2(X)}(-1))=h^{3-i}(\cO_{F_2(X)}(1)).$$
Similarly, we obtain $\chi(\cO_{F_2(X)}(2))=2816$. 
We also get
$$\chi(\cO_{F_2(X)}(3))=h^0(\cO_{F_2(X)}(3)) = 16,896 = 2^9 \cdot 3 \cdot 11,  $$
where the first equality is Kodaira vanishing.
Thus we have enough data to extract the Hilbert polynomial of $F_2(X)$
$$\chi(\cO_{F_2(X)}(m))=\frac{2^7\cdot 11}{3}(m-1) ( 5 (m-1)^2 -2).$$
Hirzebruch-Riemann-Roch allows us to compute $\chi(\Omega^1_{F_2(X)})$
as well
$$\chi(\Omega^1_{F_2(X)}) =15,616= 2^{8} \cdot 61.$$

\section{Koszul computations}
\label{sect:koszul}

\subsection{General set-up}
Let $\cE$ denote a vector bundle and $s\in \Gamma(\cE)$ a section
such that the degeneracy locus $Z=\{s=0\}$ has codimension equal
to the rank $R$ of $\cE$. Then we have a resolution
$$0 \ra \bigwedge^{R} \cE^* \ra \bigwedge^{R-1} \cE^* \ra
\cdots \ra \cE^* \ra \cO  \ra \cO_Z \ra 0$$
and the Koszul complex
$$0 \ra \bigwedge^{R} \cE^* \ra \bigwedge^{R-1} \cE^* \ra
\cdots \ra \cE^* \ra \cO  \ra 0.$$
The arrow
$$\iota_s: \bigwedge^n \cE^* \ra \bigwedge^{n-1} \cE^*$$
is contraction by $s$. 
The hypercohomology spectral sequence gives
$$E_1^{p,q}= H^{q}(\bigwedge^{-p} \cE^*) \Rightarrow H^{p+q}(\cO_Z),$$
where $p=-R,\ldots,0$.

Following \cite{Manivel}, we may use this to compute the cohomology of 
degeneracy loci over Grassmannians $\bG(n,d-1)=\Gr(n+1,V)$. 
Given a $d$-dimensional vector space $V$ and
integers $\lambda = (\lambda_1,\ldots,\lambda_k)$
with $\lambda_1 \ge \lambda_2  \ge \cdots  \ge \lambda_k$
and $\lambda_k=0$ for $k>d$, let 
$\Gamma^{\lambda} V$ denote the associated Schur functor.
We observe the convention $\Gamma^{\lambda}V=0$ for sequences
of integers $\lambda$ failing the decreasing or vanishing conditions.

We keep track of the decompositions using the dictionary between 
Schubert calculus and tensor products
$$\Gamma^{\lambda} V \otimes \Gamma^{\mu} V = 
\sum_{\nu} c_{\lambda, \mu}^{\nu} \Gamma^{\nu} V,
$$ 
where the multiplicies are defined by
$$\sigma_{\lambda}\sigma_{\mu}= \sum_{\nu} c_{\lambda,\mu}^{\nu} \sigma_{\nu}.$$

\subsection{Planes in intersections of three quadrics in $\bP^9$}
Consider the bundle
$$\cE = \Sym^2(S^*)^{\oplus 3}$$
and $s\in \Gamma(\cE)$ with $F_2(X)=\{s=0\}$. 
We have the associated Koszul resolution
$$0 \ra \bigwedge^{18} \cE^* \ra \cdots \ra \bigwedge^r \cE^* \ra
\cdots \ra \cE^* \ra \cO  \ra \cO_{F_2(X)} \ra 0.$$
The terms of the Koszul complex decompose into direct sums of products
$$\bigwedge^{e_1} \Sym^2(S) \otimes \bigwedge^{e_2} \Sym^2(S) \otimes
\bigwedge^{e_3} \Sym^2(S), \quad e_1+e_2+e_3=r.$$
The hypercohomology spectral sequence gives
$$E_1^{p,q}= H^{q}(\bigwedge^{-p} \cE^*) \Rightarrow H^{p+q}(\cO_{F_2(X)}),$$
where $p=-18,\ldots,0$.

Recall the Weyl character formula for $\operatorname{GL}(V)$
representations:
$$\dim \Gamma^{\lambda}(V) = \prod_{i<j} \frac{\lambda_i-\lambda_j+j-i}{j-i}.$$
Assume $S$ has rank three then
$$\operatorname{rank} \Gamma^{(a_1,a_2,a_3)}S = (a_1-a_2+1)(a_1-a_3+2)(a_2-a_3+1)/2.$$

We observe the dictionary
$$\begin{array}{rcl}
\Sym^2(S) & \sim & \sigma_2 \\
\bigwedge^2 \Sym^2(S) & \sim & \sigma_{3,1} \\
\bigwedge^3 \Sym^2(S) & \sim & \sigma_{3,3} + \sigma_{4,1,1} \\
\bigwedge^4 \Sym^2(S) & \sim & \sigma_{4,3,1} \\
\bigwedge^5 \Sym^2(S) & \sim & \sigma_{4,4,2} \\
\bigwedge^6 \Sym^2(S) & \sim & \sigma_{4,4,4}
\end{array}
$$
whence
\begin{align*}
\Sym^2(S) &\simeq \Gamma^{(2)}(S) \\
\bigwedge^2 \Sym^2(S) & \simeq \Gamma^{(3,1)}(S) \\
\bigwedge^3 \Sym^2(S) & \simeq \Gamma^{(3,3)}(S) \oplus \Gamma^{(4,1,1)}(S) \\
\bigwedge^4 \Sym^2(S) & \simeq \Gamma^{(4,3,1)}(S) \\
\bigwedge^5 \Sym^2(S) & \simeq \Gamma^{(4,4,2)}(S) \\
\bigwedge^6 \Sym^2(S) & \simeq \Gamma^{(4,4,4)}(S)
\end{align*}

Note also \cite[Cor.~1.3]{Jiang} that 
$$H^j(\Gr(3,10),\Gamma^{\lambda} S) = 0 $$
whenever $j$ is not divisible by $7$ and for all but
one value of $j$.  

We tabulate the representations $\Gamma^a S, a=(a_1,a_2,a_3)$
appearing in each $\bigwedge^r \cE^*$:
{\small
$$
\begin{array}{r|l}
r & \text{decomposition} \\
\hline
0 & {\color{red} \Gamma^{(0,0,0)}} \\
1 & 3\Gamma^{(2)} \\
2 & 3\Gamma^{(4)} + 6\Gamma^{(3,1)} + 3\Gamma^{(2,2)} \\
3 & \Gamma^{(6)}+ 8\Gamma^{(5,1)} + 9\Gamma^{(4,2)} + 10\Gamma^{(4,1,1)} 
  +10\Gamma^{(3,3)}+8\Gamma^{(3,2,1)}+\Gamma^{(2,2,2)}\\
4 & 3\Gamma^{(7,1)}+9\Gamma^{(6,2)}+15\Gamma^{(6,1,1)}+18\Gamma^{(5,3)}
  +24\Gamma^{(5,2,1)}+6\Gamma^{(4,4)} \\
  & +33\Gamma^{(4,3,1)}+9\Gamma^{(4,2,2)}
  +15\Gamma^{(3,3,2)} \\
5 & 3\Gamma^{(8,2)}+{\color{red} 6\Gamma^{(8,1,1)}}+9\Gamma^{(7,3)}+24\Gamma^{(7,2,1)}+18\Gamma^{(6,4)}+54\Gamma^{(6,3,1)} \\
  & +21\Gamma^{(6,2,2)}+6\Gamma^{(5,5)}+48\Gamma^{(5,4,1)}+54\Gamma^{(5,3,2)}+39\Gamma^{(4,4,2)}+30\Gamma^{(4,3,3)}\\

6& \Gamma^{(9,3)}+8\Gamma^{(9,2,1)}+8\Gamma^{(8,4)}+27\Gamma^{(8,3,1)}+19\Gamma^{(8,2,2)}+9\Gamma^{(7,5)}\\
 & + 64\Gamma^{(7,4,1)}+62\Gamma^{(7,3,2)}+10\Gamma^{(6,6)}+53\Gamma^{(6,5,1)}+117\Gamma^{(6,4,2)}+
	56\Gamma^{(6,3,3)} \\
  &+ 46\Gamma^{(5,5,2)}+88\Gamma^{(5,4,3)}+38\Gamma^{(4,4,4)}\\

7& 3\Gamma^{(10,3,1)}+6\Gamma^{(10,2,2)}+3\Gamma^{(9,5)}+24\Gamma^{(9,4,1)}+27\Gamma^{(9,3,2)}+6\Gamma^{(8,6)} \\
 & + 42\Gamma^{(8,5,1)} + 93\Gamma^{(8,4,2)} + 36\Gamma^{(8,3,3)} + 3\Gamma^{(7,7)}+48\Gamma^{(7,6,1)}+ 132\Gamma^{(7,5,2)}\\
 & +144\Gamma^{(7,4,3)}+66\Gamma^{(6,6,2)}+138\Gamma^{(6,5,3)}+114\Gamma^{(6,4,4)}+60\Gamma^{(5,5,4)} \\

8 & 3\Gamma^{(11,3,2)}+9\Gamma^{(10,5,1)}+24\Gamma^{(10,4,2)}+9\Gamma^{(10,3,3)}+3\Gamma^{(9,7)}+24\Gamma^{(9,6,1)} \\
  & +75\Gamma^{(9,5,2)}+72\Gamma^{(9,4,3)}+24\Gamma^{(8,7,1)}+102\Gamma^{(8,6,2)}+168\Gamma^{(8,5,3)}+99\Gamma^{(8,4,4)}\\
  & +69\Gamma^{(7,7,2)}+168\Gamma^{(7,6,3)}+213\Gamma^{(7,5,4)}+96\Gamma^{(6,6,4)}+75\Gamma^{(6,5,5)} \\
9 & \Gamma^{(12,3,3)}+9\Gamma^{(11,5,2)}+8\Gamma^{(11,4,3)}+9\Gamma^{(10,7,1)}+36\Gamma^{(10,6,2)} + 63\Gamma^{(10,5,3)} \\
  & +28\Gamma^{(10,4,4)}+{\color{red} \Gamma^{(9,9)}}+8\Gamma^{(9,8,1)}+63\Gamma^{(9,7,2)}+128\Gamma^{(9,6,3)}+142 \Gamma^{(9,5,4)} \\
  & +28\Gamma^{(8,8,2)}+142\Gamma^{(8,7,3)}+216\Gamma^{(8,6,4)}+146\Gamma^{(8,5,5)} + 146\Gamma^{(7,7,4)}\\
  & +160\Gamma^{(7,6,5)}+20\Gamma^{(6,6,6)}
\end{array}
$$
}
The terms in {\color{red} red} contribute to the cohomology \cite[Th.~1.2]{Jiang}. 
The representations for $r>9$ 
may be read off via duality
$$\bigwedge^{r} \cE^* = \det(\cE^*) \otimes \bigwedge^{18-r} \cE, \quad
\Gamma^{(a_1,a_2,a_3)} \times \Gamma^{(12-a_3,12-a_2,12-a_1)} \ra \Gamma^{(12,12,12)}.$$

{\small
$$
\begin{array}{r|l}
r & \text{decomposition} \\
\hline
18 & {\color{red} \Gamma^{(12,12,12)}} \\
17 & {\color{red} 3\Gamma^{(12,12,10)}} \\
16 & 3\Gamma^{(12,12,8)} + 6\Gamma^{(12,11,9)} + {\color{red} 3\Gamma^{(12,10,10)}} \\
15 & \Gamma^{(12,12,6)}+ 8\Gamma^{(12,11,7)} + 9\Gamma^{(12,10,8)} + 10\Gamma^{(11,11,8)} 
  +10\Gamma^{(12,9,9)}+8\Gamma^{(11,10,9)}\\
   & +{\color{red} \Gamma^{(10,10,10)}}\\

14 & 3\Gamma^{(12,11,5)}+9\Gamma^{(12,10,6)}+15\Gamma^{(11,11,6)}+18\Gamma^{(12,9,7)}
  +24\Gamma^{(11,10,7)}+6\Gamma^{(12,8,8)} \\
  & +33\Gamma^{(11,9,8)}+9\Gamma^{(10,10,8)}
  +15\Gamma^{(10,9,9)} \\

13 & 3\Gamma^{(12,10,4)}+6\Gamma^{(11,11,4)}+9\Gamma^{(12,9,5)}+24\Gamma^{(11,10,5)}+18\Gamma^{(12,8,6)}+54\Gamma^{(11,9,6)} \\
  & +21\Gamma^{(10,10,6)}+6\Gamma^{(12,7,7)}+48\Gamma^{(11,8,7)}+54\Gamma^{(10,9,7)}+39\Gamma^{(10,8,8)}+30\Gamma^{(9,9,8)}\\

12& \Gamma^{(12,9,3)}+8\Gamma^{(11,10,3)}+8\Gamma^{(12,8,4)}+27\Gamma^{(11,9,4)}+19\Gamma^{(10,10,4)}+9\Gamma^{(12,7,5)}\\
 & + 64\Gamma^{(11,8,5)}+62\Gamma^{(10,9,5)}+10\Gamma^{(12,6,6)}+53\Gamma^{(11,7,6)}+117\Gamma^{(10,8,6)}\\
  &+56\Gamma^{(9,9,6)} 
  + 46\Gamma^{(10,7,7)}+88\Gamma^{(9,8,7)}+38\Gamma^{(8,8,8)}\\

11&{\color{red} 3\Gamma^{(11,9,2)}}+{\color{red} 6\Gamma^{(10,10,2)}}+3\Gamma^{(12,7,3)}+24\Gamma^{(11,8,3)}+27\Gamma^{(10,9,3)}+6\Gamma^{(12,6,4)} \\
 & + 42\Gamma^{(11,7,4)} + 93\Gamma^{(10,8,4)} + 36\Gamma^{(9,9,4)} + 3\Gamma^{(12,5,5)}+48\Gamma^{(11,6,5)}+ 132\Gamma^{(10,7,5)}\\
 & +144\Gamma^{(9,8,5)}+66\Gamma^{(10,6,6)}+138\Gamma^{(9,7,6)}+114\Gamma^{(8,8,6)}+60\Gamma^{(8,7,7)} \\

10 & {\color{red} 3\Gamma^{(10,9,1)}}+9\Gamma^{(11,7,2)}+24\Gamma^{(10,8,2)}+{\color{red} 9\Gamma^{(9,9,2)}}+3\Gamma^{(12,5,3)}+24\Gamma^{(11,6,3)} \\
  & +75\Gamma^{(10,7,3)}+72\Gamma^{(9,8,3)}+24\Gamma^{(11,5,4)}+102\Gamma^{(10,6,4)}+168\Gamma^{(9,7,4)}\\
   &+99\Gamma^{(8,8,4)}
 +69\Gamma^{(10,5,5)}+168\Gamma^{(9,6,5)}+213\Gamma^{(8,7,5)}+96\Gamma^{(8,6,6)}+75\Gamma^{(7,7,6)} 
\end{array}
$$
}

And then we record those contributing cohomology in degree $j$,
using  \cite[Th.~1.2]{Jiang}. Note that the sequence
$$(-1,-2,-3,-4,-5,-6,-7,a_1-8,a_2-9,a_3-10)$$
must have no repeating integers as these yield `singular'
weights. Thus we must have $a_1\ge 8$; if $a_1=8$ then $a_2=1,0$, etc.
$$\begin{array}{r|l}
j & \text{weights $a$ contributing } \\
\hline \hline
0 & (0,0,0) \\
\hline
7  & \text{no contributions in } \bigwedge^4 \cE^* \\
   & (8,1,1) \\
   & \text{no contributions in } \bigwedge^6 \cE^* \\
\hline
14 & (9,9,0) \\
   & (10,9,1), (9,9,2) \\
   & (10,10,2), (11,9,2) \\
   &  \text{no contributions in } \bigwedge^{12} \cE^* \\
\hline
21 & (10,10,10)  \\
   & (12,10,10) \\
  & (12,12,10)  \\
 & (12,12,12)
\end{array}$$

The degree $7$ cohomology contributes through
$$H^7(\bigwedge^5 \cE^*)=H^7((\Gamma^{(8,1,1)}S)^{\oplus 6})
\simeq (\Gamma^{(1,1,1,1,1,1,1,1,1,1)}V)^{\oplus 6}
\simeq (\det(V))^{\oplus 6} \simeq \bC^6.$$
Here $V$ is the standard $10$-dimensional representation.

\begin{prop} \label{prop:getsix}
We have
$$H^0(\cO_{F_2(X)})=\bC, \quad H^1(\cO_{F_2(X)})=0.$$
For higher degrees, we have
$$\bC^6 \simeq
E_1^{-5,7}=E_2^{-5,7}=\cdots = E_{\infty}^{-5,7}=H^2(\cO_{F_2(X)}).$$
Consequently, we deduce that 
$$h^3(\cO_{F_2(X)})=2816 + 1 + 6 - 0 = 2823.$$
\end{prop}
\begin{proof} The only term in degree zero is $E_1^{(0,0)}$ and there are no terms in degree one, which gives the degree $0$ and $1$ cohomology; see also Prop.~\ref{prop:DebMan} and \cite{DebMan}. The spectral sequence is supported at the following values of $(p,q)$:
\begin{align*}
(0,0), (-5,7), (-9,14),& (-10,14),(-11,14), \\
 (-15,21), (-16,21),&(-17,21), (-18,21).
 \end{align*}
All the values after $(-5,7)$ have degree $p+q \ge 3$ and $p\le -5$, thus 
do not receive arrows from $E_r^{-5,7}$. And clearly there are no maps from 
the degree $(0,0)$ term, abutting to $\Gamma(\cO_{F_2(X)})$. This yields
the equalities asserted above.
\end{proof}

The degree $14$ cohomology contributes through
\begin{eqnarray*}
H^{14}(\bigwedge^{11} \cE^*) &=& H^{14}((\Gamma^{(10,10,2)}S)^{\oplus 6})\oplus
					(\Gamma^{(11,9,2)}S)^{\oplus 3})) \\
			     &=& (\Gamma^{(3,3,2,2,2,2,2,2,2,2)}V)^{\oplus 6}
			\oplus (\Gamma^{(4,2,2,2,2,2,2,2,2)}V)^{\oplus 3} \\
		             &\simeq & (\bC^{45})^{6} \oplus (\bC^{55})^{3} 
					\simeq \bC^{435} \\
H^{14}(\bigwedge^{10} \cE^*) &=& H^{14}((\Gamma^{(9,9,2)}S)^{\oplus 9})\oplus
					(\Gamma^{(10,9,1)}S)^{\oplus 3})) \\
			   &=& (\Gamma^{(2,2,2,2,2,2,2,2,2,2)}V)^{\oplus 9}
			\oplus (\Gamma^{(3,2,2,2,2,2,2,2,2,1)}V)^{\oplus 3} \\
			&\simeq & (\bC)^9 \oplus (\bC^{99})^3 \simeq  \bC^{306} \\
H^{14}(\bigwedge^9 \cE^*) &=&  H^{14}(\Gamma^{(9,9)}S)  \\
 			  &=& \Gamma^{(2,2,2,2,2,2,2,2,2,0)}V
					\simeq \bC^{55} 
\end{eqnarray*}

The degree $21$ cohomology contributes through:
\begin{eqnarray*}
H^{21}(\bigwedge^{18} \cE^*) &=& H^{21}(\Gamma^{12,12,12}S) \\
		&=&\Gamma^{(5,5,5,3,3,3,3,3,3,3)}V\simeq \bC^{4950} \\
H^{21}(\bigwedge^{17} \cE^*) &=& H^{21}((\Gamma^{12,12,10}S)^{\oplus 3}) \\
		&=&(\Gamma^{(5,5,3,3,3,3,3,3,3,3)}V)^{\oplus 3}
			\simeq (\bC^{825})^3 \simeq \bC^{2475} \\
H^{21}(\bigwedge^{16} \cE^*) &=& H^{21}((\Gamma^{12,10,10}S)^{\oplus 3}) \\
		&=&(\Gamma^{(5,3,3,3,3,3,3,3,3,3)}V)^{\oplus 3} \simeq (\bC^{55})^3 \simeq \bC^{165}  \\
H^{21}(\bigwedge^{15} \cE^*) &=& H^{21}(\Gamma^{10,10,10}S) \\
		&=&\Gamma^{(3,3,3,3,3,3,3,3,3)}V \simeq \bC
\end{eqnarray*}

\begin{rema}
Standard properties of spectral sequences imply
$$-2816=\chi(\cO_{F_2(X)})= \sum (-1)^{p+q}
H^{q}(\bigwedge^{-p} \cE^*)= \sum (-1)^{a+p} H^{7a}(\bigwedge^{-p} \cE^*).$$
We compute 
$$-2816 =   1+6+(-55+306-435)+(-4950+2475-165+1) ,$$
thus everything checks.
\end{rema}

\begin{ques} The simplest structure on the spectral sequence would be to have exactness 
in all degrees where there is no contribution to cohomology, i.e., degeneration at $E_2$. Thus the only
nonzero terms in the second page would be
$$E^{-18,21}_2 \simeq \bC^{4950-2475+165-1} = \bC^{2639}, \quad
E^{-11,14}_2 \simeq \bC^{435-306+55}  = \bC^{184},$$
as well as
$$E^{-5,7}_2 \simeq \bC^6, \quad E^{0,0}_2\simeq \bC.$$
Does this degeneration occur? It would imply a nontrivial filtration on the holomorphic three-forms
of $F_2(X)$.  
\end{ques}

\begin{prop}
The Hodge diamond of $F_2(X)$ is
$$\begin{array}{ccccccc}
  &   &   &  1  &  &  &  \\
   &   &   0  &  & 0 & & \\
   &  6 &  & 62 & & 6 & \\
 2823  & &15684 & & 15684 & & 2823
 \end{array}  
$$
\end{prop}

\begin{proof}
We recall that 
\begin{itemize}
\item{The Picard rank $\rho(F_2(X))=1$
when $X$ is very general \cite[Th.~0.3]{Jiang}.}
\item{Let $\Sigma\ra \bP^2$ denote the double cover branched along degeneracy curve;
there exists an isometry on primitive cohomology
$$H^6(X,\bZ)_{\prim} \hookrightarrow H^2(\Sigma,\bZ)(-2)_{\prim},$$ 
realizing the former as an index-two sublattice of the latter \cite[Th.~0.1]{ogrady86}.}
\end{itemize}
The latter observation gives the Hodge numbers of $X$ displayed in
Section~\ref{subsect:recall}.

The incidence correspondence 
$$\begin{array}{rcccl}
	&	& Z & &  \\
	 &\stackrel{f}{\swarrow} & & \stackrel{g}{\searrow} & \\
X  &     &   & & F_2(X)
\end{array}
$$
induces Abel-Jacobi maps
$$
\alpha_1:g_*f^*:H^6(X,\bZ) \ra H^2(F_2(X),\bZ)(-2), 
$$
$$
\alpha_2:f_*g^*:H^4(F_2(X),\bZ) \ra H^6(X,\bZ)(1).
$$
Letting 
$$
L:H^2(F_2(X),\bZ) \ra H^4(F_2(X),\bZ)
$$ denote intersection by the
hyperplane class $\sigma_1$,
the composition
$$\alpha_2 \circ L \circ \alpha_1:H^6(X,\bZ) \ra H^6(X,\bZ)$$
forces the cohomology of $X$ and $F_2(X)$ to be tightly
intertwined. Such constructions are used to establish
Grothendieck's version of the Hodge conjecture for 
varieties with sparse Hodge diamond; see for example
\cite{VoisinConiveau}.

For our immediate purpose, we can apply the Main Theorem of
\cite{Shimada} to conclude that the cylinder map $\alpha_1$
injects the primitive cohomology $H^6(X,\bZ)_{\prim}$ into 
$H^2(F_2(X),\bZ)$. Since we computed 
$$h^{0,2}(F_2(X))=6$$
in Proposition~\ref{prop:getsix}, we conclude that 
$$
H^2(F_2(X),\bQ)_{\prim} \simeq H^6(X,\bQ)(2)_{\prim} \oplus \bQ(-1)^N
$$
for some $N$.
The last summand is of Hodge-Tate type. 
However, $N=0$ by Jiang's result on the Picard group.
\end{proof}

\bibliographystyle{alpha}
\bibliography{3quad}
\end{document}